A box in a college bookstore contains books, and each book in the box is a history book, an English Book or a Science book. If 1/3 of these books are History books and 1/6 are English books, What fraction of the books are science books?
Answer by rfer(16322)

2/6+1/6=3/6=1/2 H&E
Science=1/2